OpenStudy (anonymous):

complete the square using this equation x^2-4.7x=-2.8

OpenStudy (anonymous):

(x - 2.35)^2 - 2.35^2 = -2.8 (x - 2.35)^2 = -2.8 + 2.35^2 = 2.7225 x - 2.35 = +-sqrt 2.7225 x = sqrt 2.7225 + 2.35 = 1.65 + 2.35 = 4 or -sqrt 2.7225 + 2.35 = -1.65 + 2.35 = 0.70